OpenStudy (anonymous):

i need to create a list of 10 "reasonings" that you can use in proofs. can someone please help! im horrible with proofs!

OpenStudy (anonymous):

well when you do proofs what do you usually use?

OpenStudy (anonymous):

p _ _ p _ _ t i e s

OpenStudy (anonymous):

=P

OpenStudy (anonymous):

i cant do proofs because i dont understand them. i need to make a list of 10 and include what they look likeD:

OpenStudy (anonymous):

when you do proofs you usually use properties

OpenStudy (anonymous):

or identities

OpenStudy (anonymous):

such as the additive identity, commutative identity,

OpenStudy (anonymous):

i can give you a list of them ane it's funny because there is 10 identities to prove a space

OpenStudy (anonymous):

would you be kind enough to show what they look like also?(:

OpenStudy (anonymous):

i have a 29% in geometry, amd you can see why....

OpenStudy (anonymous):

well usually he question is like Prove that 1=1 or something and you use identities to solve all the way to 1=1

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

something like this there is a statement such as "show that this (equation) = this solution

OpenStudy (anonymous):

that doesnt look like our proofs

OpenStudy (anonymous):

then you use identities to solve it

OpenStudy (anonymous):

do you have a picture of one of your's i can do one and show you what it'd look like

OpenStudy (anonymous):

like one from your book or somethin

OpenStudy (anonymous):

im cyberschooled.

OpenStudy (anonymous):

like its just simple proofs

OpenStudy (anonymous):

alright hold on do you have a question of one or something lol you say it doesn't look like one you've seen so can i see one you've seen

OpenStudy (anonymous):

like it looks like it, just not that compicated!

OpenStudy (anonymous):

id get one but they are downstairs and i have 2 puppies down there and i dont want to wake them up. haha

OpenStudy (anonymous):

hmm let me see if i have one in my old book

OpenStudy (anonymous):

we use ones that have like a diagram of a triangle and has a given and all

OpenStudy (experimentx):

smells like geometry

OpenStudy (experimentx):

do you have some dummy question??

OpenStudy (anonymous):

nno

OpenStudy (anonymous):

Additive identity: A + B = B + A

OpenStudy (anonymous):

Commutative: a*b=b*a

OpenStudy (experimentx):

Assuming straight angle is 180